# Leveraging Modality Tags for Enhanced Cross-Modal Video Retrieval

### v1

@@ V Qualitative Results
Fig. 6 shows qualitative results on all datasets. We compare retrieved ranks of MAC-VR against our baseline without visual and textual tags. In general, both visual and textual tags add complementary information extracted from the video and the caption that directly help retrieval. For example, consider the query a class is being introduced to a digital reading device and its video top right of Fig. 6. Visual tags, student, technology, learning, and textual tags, initial setup, introduction, training, add additional information to describe the action shown in the video/described in the caption. We acknowledge there are cases where tags are harmful. For example, given the query a person is cooking on stage we can see that singing is one of the visual tags extracted. The VLM extracted this tag from the last frame where there is a person singing on stage as the scene changed. Additional results are shown in Sec. G in appx.
We align the visual and textual concepts with the corresponding auxiliary modality-specific concepts by introducing our Alignment Loss $\mathcal{L_{A}}$ (Sec. III-C). Fig. 7 shows the t-SNE of visual and textual concepts with/without the auxiliary tags on MSR-VTT and Charades ($1000$ samples), see Sec. F in the appx. for additional t-SNE plots. Our method better groups concepts into clusters over the baseline.
@@
Discussion and Future Work. We find that modality-specific tags are certainly beneficial for video retrieval, but also discover that they can be harmful for specific queries. This could represent either correct tags that are redundant or incorrect tags due to errors in tag extraction. Foundation models, i.e. VLMs and LLMs, tend to hallucinate the content of the output meaning that the generated content might stray from factual reality or include fabricated information [99, 100]. Finally, we treat all the visual and textual tags with the same importance, it is possible that some generated words can be more or less discriminative than others based on uniqueness and/or relatedness to the video/caption. We leave this exploration for future work.
VI Conclusion
@@
Acknowledgements
This work used public datasets and was supported by EPSRC UMPIRE (EP/T004991/1) and EPSRC Program Grant Visual AI (EP/T028572/1).
Appendix A Tag Extraction
In Sec. III-B, we have described how we extracted tags from a video and its corresponding caption, here we provide additional information and the prompts used to extract tags for the video and text modalities. For all the datasets except DiDeMo, we use the video and its corresponding caption to generate visual and textual tags, respectively. For DiDeMo, due to it being used for video-paragraph retrieval, we first concatenate the captions from the video to form the associated paragraph. Then, we generate tags from the entire video and paragraph.
The prompt used as input of VideoLLaMA2 [37] to extract visual tags is: {mdframed}[backgroundcolor=blue!10, linecolor=red, linewidth=2pt, roundcorner=10pt, shadow=true]
A general tag of an action is a fundamental and overarching idea that encapsulates the essential principles, commonalities, or recurrent patterns within a specific behavior or activity, providing a higher-level understanding of the underlying themes and purpose associated with that action.
@@ Appendix A Tag Extraction
The prompt for Llama3.1 [95] to extract textual tags is: {mdframed}[backgroundcolor=blue!10, linecolor=red, linewidth=2pt, roundcorner=10pt, shadow=true] A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions.
@@
USER: You are a conversational AI agent. You typically extract general tags of an action.
@@
A general tag of an action is a fundamental and overarching idea that encapsulates the essential principles, commonalities, or recurrent patterns within a specific behavior or activity, providing a higher-level understanding of the underlying themes and purpose associated with that action.
Given the following action: 1) {}
What are the top 10 general tags of the above action? Use one or two significant words per tag and do not give any explanation.
ASSISTANT:
@@
Note that {} will be replaced with the caption. We did not provide any examples in the prompts (i.e. in context learning) as we found that this led to the foundation models hallucinating these examples in the output. Rather, we found that the foundation models were able to generate reasonable outputs for both video and text. We do not use any strategy to avoid the foundation models hallucinating as spot-checking the results found them to be clean enough for our purposes. The only post-processing strategy we adopted was to clean the output of the models in order to get the corresponding tags: we remove punctuation; stopwords; extracted tags that contain a noun and a verb to avoid the presence of complete sentences as tags; and tags larger than $3$ words. In Fig. 8 we show additional examples of tags for all the datasets.

### v3

@@ 5 Qualitative Results
Fig. 6 shows qualitative results across all the datasets. We compare retrieved ranks of MAC-VR against our baseline without visual and textual tags. In general, both visual and textual tags add complementary information extracted from the video and the caption that directly help retrieval. For example, consider the query a class is being introduced to a digital reading device and its video top right of Fig. 6. Visual tags, student, technology, learning, and textual tags, initial setup, introduction, training, add additional information to describe the action shown in the video/described in the caption. We acknowledge there are cases where tags are harmful. Additional results, including failure cases are shown in Sec. J in supp.
Figure 6: Qual. results across all datasets comparing baseline/MAC-VR resp. left/right rank.
Figure 7: t-SNE plot of visual and textual concepts on MSR-VTT-9k without (left) and with using auxiliary modality-specific tags (right).
@@
Fig. 7 shows the t-SNE plot of MAC-VR of visual and textual concepts with/without auxiliary modality-specific tags on MSR-VTT-9k. We can see that the use of auxiliary modality-specific tags and our Alignment Loss $\mathcal{L_{A}}$ helps to better distinguish the different concepts and have better clusters in the t-SNE plot. Additional t-SNE plots are shown in the supp. Discussion and Future Work. We find that modality-specific tags are helpful for video retrieval, but they can be harmful for specific queries. This could represent either correct but redundant tags or incorrectly extracted tags. Foundation models, i.e. VLMs and LLMs, often hallucinate the content of the output, generating factual reality or including fabricated information [Rawte et al.(2023)Rawte, Sheth, and Das, Sahoo et al.(2024)Sahoo, Meharia, Ghosh, Saha, Jain, and Chadha]. Finally, we treat all the visual/textual tags with the same importance; it is possible that some words can be more or less discriminative than others based on uniqueness and/or relatedness to the video/caption. We leave this exploration for future work.
6 Conclusion
@@
Acknowledgements This work used public datasets and was supported by EPSRC Fellowship UMPIRE (EP/T004991/1) and EPSRC Program Grant Visual AI (EP/T028572/1).
Appendix A Supplementary Contents
In the supplementary, we present further information about MAC-VR and further ablations over our design choices. In Sec. B we introduce the Querybank Normalisation (QB) [Bogolin et al.(2022)Bogolin, Croitoru, Jin, Liu, and Albanie] and Dual Softmax (DSL) [Cheng et al.(2021)Cheng, Lin, Wu, Yang, and Shen]. Then, Sec. C provides additional details on the process of extracting tags from the foundation models, as well as further information about the extracted tags. In Sec. D, we describe in detail all the datasets used in the main paper. Then, in Sec. E, we present the full tables shown in the main paper. Then, in Sec. F, we ablate additional architecture details of MAC-VR. After, we showcase the effect of using tags extracted from different foundation models in Sec. G. Then, we present a comparison to auxiliary captions in Sec. H, further t-SNE plots of MAC-VR in Sec. I, and additional qualitative results in Sec. J.
Appendix B Inference Strategies
The QB strategy [Bogolin et al.(2022)Bogolin, Croitoru, Jin, Liu, and Albanie] was introduced to mitigate the hubness problem of high-dimensional embedding spaces [Radovanovic et al.(2010)Radovanovic, Nanopoulos, and Ivanovic], where a small subset of samples tends to appear far more frequently among the k-nearest neighbours queries. To mitigate this, the similarities between embeddings are altered to minimise the influence of hubs. A set of samples (the querybank) is sampled from the queries within the training set and used as a probe to measure the hubness of the gallery at test time. More formally, given the vector of unnormalised similarities at test time, $S(v_{j},t_{i})$ and a probe matrix $P$, in which each row is a probe vector of similarities between the querybank and each element in the gallery, we re-weight the query-gallery similarities: $\eta_{q}=\text{QB}(S(v_{j},t_{i}),P)$, where the QB is the querybank normalisation function using Dynamic Inverted Softmax (DIS), introduced in [Bogolin et al.(2022)Bogolin, Croitoru, Jin, Liu, and Albanie].
The second commonly used inference strategy, the DSL strategy [Cheng et al.(2021)Cheng, Lin, Wu, Yang, and Shen], introduces an intrinsic prior of each pair in a batch to correct the similarity matrix and achieves the dual optimal match. In practice, we modify the original $S(v_{j},t_{i})$ by multiplying it with a prior $r_{i,j}$. Therefore, we can define the new similarity matrix as $\hat{S}(v_{j},t_{i})=r_{i,j}S(v_{j},t_{i})$, where the prior is defined as $r_{i,j}=\frac{exp(\tau_{r}S(v_{i},t_{i}))}{\sum_{j}exp(\tau_{r}S(v_{i},t_{j}))}$, where $\tau_{r}$ is a temperature hyper-parameter to smooth the gradients. While this strategy can be used both in training and inference, it is now regularly used only during inference.
Appendix C Tag Extraction
In Sec. 3.2, we have described how we extracted tags from a video and its corresponding caption, here we provide additional information and the prompts used to extract tags for the video and text modalities. For all the datasets except DiDeMo, we use the video and its corresponding caption to generate visual and textual tags, respectively. For DiDeMo, due to it being used for video-paragraph retrieval, we first concatenate the captions from the video to form the associated paragraph. Then, we generate tags from the entire video and paragraph.
The prompt used as input of VideoLLaMA2 [Cheng et al.(2024b)Cheng, Leng, Zhang, Xin, Li, Chen, Zhu, Zhang, Luo, Zhao, et al.] to extract visual tags is: A general tag of an action is a fundamental and overarching idea that encapsulates the essential principles, commonalities, or recurrent patterns within a specific behaviour or activity, providing a higher-level understanding of the underlying themes and purpose associated with that action.
@@ Appendix C Tag Extraction
The prompt for Llama3.1 [Dubey et al.(2024)Dubey, Jauhri, Pandey, Kadian, Al-Dahle, Letman, Mathur, Schelten, Yang, Fan, Goyal, Hartshorn, Yang, Mitra, Sravankumar, Korenev, Hinsvark, Rao, Zhang, Rodriguez, Gregerson, Spataru, Rozière, Biron, Tang, Chern, Caucheteux, Nayak, Bi, Marra, McConnell, Keller, Touret, Wu, Wong, Ferrer, Nikolaidis, Allonsius, Song, Pintz, Livshits, Esiobu, Choudhary, Mahajan, Garcia-Olano, Perino, Hupkes, Lakomkin, AlBadawy, Lobanova, Dinan, Smith, Radenovic, Zhang, Synnaeve, Lee, Anderson, Nail, Mialon, Pang, Cucurell, Nguyen, Korevaar, Xu, Touvron, Zarov, Ibarra, Kloumann, Misra, Evtimov, Copet, Lee, Geffert, Vranes, Park, Mahadeokar, Shah, van der Linde, Billock, Hong, Lee, Fu, Chi, Huang, Liu, Wang, Yu, Bitton, Spisak, Park, Rocca, Johnstun, Saxe, Jia, Alwala, Upasani, Plawiak, Li, Heafield, Stone, and et al.] to extract textual tags is: A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions.
@@
USER: You are a conversational AI agent. You typically extract the general tags of an action.
@@
A general tag of an action is a fundamental and overarching idea that encapsulates the essential principles, commonalities, or recurrent patterns within a specific behaviour or activity, providing a higher-level understanding of the underlying themes and purpose associated with that action. Given the following action: 1) {} What are the top 10 general tags of the above action? Use one or two significant words per tag and do not give any explanation. ASSISTANT:
@@
Note that {} will be replaced with the caption. We did not provide any examples in the prompts (i.e. in context learning) as we found that this led to the foundation models hallucinating these examples in the output. Rather, we found that the foundation models were able to generate reasonable outputs for both video and text.
Figure 8: Additional examples of visual and textual tags across our datasets.
We do not use any strategy to avoid the foundation models hallucinating as spot-checking the results found them to be clean enough for our purposes. The only post-processing strategy we adopted was to clean the output of the models in order to get the corresponding tags: we remove punctuation; stopwords; extracted tags that contain a noun and a verb to avoid the presence of complete sentences as tags; and tags larger than $3$ words. Figure 8 illustrates additional tag examples for each dataset, while Table 8 summarises the number of unique visual and textual tags across all datasets.
#Unique Visual Tags (VT) #Unique Textual Tags (TT)
Datasets #Videos # Captions Train Val Test Train Val Test
MSR-VTT-9k [Gabeur et al.(2020)Gabeur, Sun, Alahari, and Schmid] 10,000 190,000 63,383 - 12,118 320,351 - 8,326
MSR-VTT-7k [Miech et al.(2019)Miech, Zhukov, Alayrac, Tapaswi, Laptev, and Sivic] 8,010 141,200 52,309 - 12,118 270,367 - 8,326
DiDeMo [Hendricks et al.(2017)Hendricks, Wang, Shechtman, Sivic, Darrell, and Russell] 10,642 10,642 50,712 10,924 10,636 34,662 9,234 8,266
TGIF [Li et al.(2016)Li, Song, Cao, Tetreault, Goldberg, Jaimes, and Luo] 100,551 100,551 105,895 28,500 29,538 176,353 44,053 45,417
Charades [Sigurdsson et al.(2016)Sigurdsson, Varol, Wang, Farhadi, Laptev, and Gupta] 9,848 9,848 41,538 - 15,360 25,406 - 9,874
YouCook2 [Zhou et al.(2018)Zhou, Xu, and Corso] 13,829 13,829 33,786 - 16,457 26,035 - 12,227
Table 8: Statistical analysis of tags after extraction.
Appendix D Datasets
MSR-VTT [Xu et al.(2016)Xu, Mei, Yao, and Rui] is commonly utilised for video retrieval. It comprises of 10K YouTube videos, each with $20$ captions. We utilise two commonly used training splits, with the same test set of 1K videos: the 9k-Train split [Gabeur et al.(2020)Gabeur, Sun, Alahari, and Schmid] and the 7k-Train split [Miech et al.(2019)Miech, Zhukov, Alayrac, Tapaswi, Laptev, and Sivic]. DiDeMo [Hendricks et al.(2017)Hendricks, Wang, Shechtman, Sivic, Darrell, and Russell] collects 10K Flickr videos annotated with paragraph captions. This dataset is evaluated in a video-paragraph retrieval manner [Luo et al.(2022b)Luo, Ji, Zhong, Chen, Lei, Duan, and Li]. TGIF [Li et al.(2016)Li, Song, Cao, Tetreault, Goldberg, Jaimes, and Luo] is a large dataset that contains more than 100K animated short GIFs collected from Tumblr, and natural language sentences annotated via crowdsourcing. We use the split proposed in [Li et al.(2019)Li, Xu, Yang, Chen, and Dong]. Charades [Sigurdsson et al.(2016)Sigurdsson, Varol, Wang, Farhadi, Laptev, and Gupta] contains $9,848$ videos of $157$ action classes in total, where each video is associated with a caption. This dataset consists of fine-grained videos of recorded human actions. We use the standard training and test splits [Lin et al.(2022)Lin, Lei, Bansal, and Bertasius]. YouCook2 [Zhou et al.(2018)Zhou, Xu, and Corso] is a fine-grained dataset that contains 14K YouTube cooking videos that cover $89$ recipes. A textual sentence describes each video clip. Following [Miech et al.(2019)Miech, Zhukov, Alayrac, Tapaswi, Laptev, and Sivic], we evaluate this dataset on the validation set.
DiCoSA* [Jin et al.(2023b)Jin, Li, Cheng, Huang, Wang, Yuan, Liu, and Chen] - IJCAI’23 47.2 73.5 83.0 2 12.9 203.7 44.4 72.6 81.2 2 14.7 198.2 41.2 71.3 81.3 2 15.9 193.8
DiCoSA* [Jin et al.(2023b)Jin, Li, Cheng, Huang, Wang, Yuan, Liu, and Chen] QB IJCAI’23 48.0 74.6 84.3 2 12.9 206.9 44.7 73.3 81.3 2 14.3 199.3 43.7 73.2 81.7 2 16.8 198.6
DiCoSA* [Jin et al.(2023b)Jin, Li, Cheng, Huang, Wang, Yuan, Liu, and Chen] DSL IJCAI’23 52.1 77.3 85.9 1 12.9 215.3 49.2 73.3 82.8 2 12.6 205.3 47.3 75.7 83.8 2 14.2 206.8
DiCoSA* [Jin et al.(2023b)Jin, Li, Cheng, Huang, Wang, Yuan, Liu, and Chen] - IJCAI’23 14.0 28.5 36.1 30 296.7 78.6 16.7 35.9 47.1 12 61.7 99.7 11.9 29.2 40.3 18 117.9 81.4
DiCoSA* [Jin et al.(2023b)Jin, Li, Cheng, Huang, Wang, Yuan, Liu, and Chen] QB IJCAI’23 14.1 28.4 36.4 30 313.6 78.9 18.8 39.5 49.0 11 60.8 107.3 12.9 31.5 42.3 16 108.7 86.7
DiCoSA* [Jin et al.(2023b)Jin, Li, Cheng, Huang, Wang, Yuan, Liu, and Chen] DSL IJCAI’23 15.5 30.1 37.8 28 290.6 83.4 18.2 39.2 49.1 11 58.9 106.5 12.8 32.2 43.0 15 103.8 88.0
Table 9: Comparison with baseline trained by using same training parameters of MAC-VR. our reproduced results with same training parameters. IS: Inference Strategy. BSBS: Batch Size. $N_{v}$: Number of Frames. The highest performance in each block is marked in bold.
MSR-VTT-9k MSR-VTT-7k DiDeMo
CLIP4Clip [Luo et al.(2022a)Luo, Ji, Zhong, Chen, Lei, Duan, and Li] - Neurocomp.’22 44.5 71.4 81.6 2 15.3 197.5 42.1 71.9 81.4 2 16.2 195.4 43.4 70.2 80.6 2 17.5 194.2
CenterCLIP [Zhao et al.(2022)Zhao, Zhu, Wang, and Yang] - SIGIR’22 44.2 71.6 82.1 2 15.1 197.9 43.7 71.3 80.8 2 16.9 195.8 - - - - - -
X-Pool [Gorti et al.(2022)Gorti, Vouitsis, Ma, Golestan, Volkovs, Garg, and Yu] - CVPR’22 46.9 72.8 82.2 2 14.3 201.9 43.9 72.5 82.3 2 14.6 198.7 - - - - - -
TS2-Net [Liu et al.(2022)Liu, Xiong, Xu, Cao, and Jin] - ECCV’22 47.0 74.5 83.8 2 13.0 205.3 - - - - - - 41.8 71.6 82.0 2 14.8 195.4
EMCL-Net [Jin et al.(2022)Jin, Huang, Liu, Wu, Ge, Song, Clifton, and Chen] - NeurIPS’22 46.8 73.1 83.1 2 - 203.0 - - - - - - 46.8 74.3 83.1 2 12.3 204.2
EMCL-Net [Jin et al.(2022)Jin, Huang, Liu, Wu, Ge, Song, Clifton, and Chen] - NeurIPS’22 - - - - - - 45.2 71.4 81.2 2 15.0 197.8 44.1 71.7 80.3 2 15.1 197.8
VoP [Huang et al.(2023)Huang, Gong, Pan, Jiang, Lv, Li, and Wang] - CVPR’23 44.6 69.9 80.3 2 16.3 194.8 42.7 68.2 79.3 2 15.9 190.2 46.4 71.9 81.5 2 13.6 199.8
MuMUR [Madasu et al.(2023)Madasu, Aflalo, Stan, Rosenman, Tseng, Bertasius, and Lal] - Inf. Retr. J.’23 46.4 72.6 82.2 2 13.9 201.2 44.8 72.0 82.5 2 - 199.3 44.4 74.3 83.1 2 - 201.8
PiDRo [Guan et al.(2023)Guan, Pei, Shao, Liu, Li, Gu, Xu, Xu, Yan, and Lam] - ICCV’23 48.2 74.9 83.3 2 12.6 206.4 - - - - - - 48.6 75.9 84.4 2 11.8 208.9
HBI [Jin et al.(2023a)Jin, Huang, Xiong, Tian, Liu, Ji, Yuan, and Chen] - CVPR’23 48.6 74.6 83.4 2 12.0 206.6 - - - - - - 46.9 74.9 82.7 2 12.1 204.5
DiffusionRet [Jin et al.(2023c)Jin, Li, Cheng, Li, Ji, Liu, Yuan, and Chen] - ICCV’23 49.0 75.2 82.7 2 12.1 206.9 - - - - - - 46.7 74.7 82.7 2 14.3 204.1
Prompt Switch [Deng et al.(2023)Deng, Chen, Qin, Chen, and Wu] - ICCV’23 47.8 73.9 82.2 - 14.4 203.9 - - - - - - - - - - - -
Cap4Video [Wu et al.(2023a)Wu, Luo, Fang, Wang, and Ouyang] - CVPR’23 49.3 74.3 83.8 2 12.0 207.4 - - - - - - 52.0 79.4 87.5 1 10.5 218.9
UCoFiA [Wang et al.(2023b)Wang, Sung, Cheng, Bertasius, and Bansal] - ICCV’23 49.4 72.1 83.5 2 12.9 205.0 - - - - - - 46.5 74.8 84.4 2 13.4 205.7
UCoFiA [Wang et al.(2023b)Wang, Sung, Cheng, Bertasius, and Bansal] - ICCV’23 - - - - - - 45.2 69.6 79.4 2 15.7 194.2 42.1 69.2 79.1 2 16.3 190.4
PAU [Li et al.(2023a)Li, Song, Gao, Zhu, and Shen] - NeurIPS’23 48.5 72.7 82.5 2 14.0 203.7 - - - - - - 48.6 76.0 84.5 2 12.9 209.1
TABLE [Chen et al.(2023b)Chen, Wang, Lin, Qi, Ma, and Shan] - AAAI’23 47.1 74.3 82.9 2 13.4 204.3 - - - - - - 47.9 74.0 82.1 2 14.3 204.0
UATVR [Fang et al.(2023)Fang, Wu, Liu, Zhou, Song, Wang, Shu, Ji, and Wang] - ICCV’23 47.5 73.9 83.5 2 12.3 204.9 - - - - - - 43.1 71.8 82.3 2 15.1 197.2
TeachCLIP [Tian et al.(2024)Tian, Zhao, Xin, Lan, and Li] - CVPR’24 46.8 74.3 - - - - - - - - - - 43.7 71.2 - - - -
MV-Adapter [Jin et al.(2024)Jin, Zhang, Gong, Xu, Deng, Wang, Zhang, Shen, and Feng] - CVPR’24 46.2 73.2 82.7 - - 202.1 - - - - - - 44.3 72.1 80.5 - - 196.9
BiC-Net [Han et al.(2024)Han, Zeng, Shi, Xiao, Chen, and Chen] - TOMM’24 39.4 75.5 86.7 2 - 201.6 32.8 68.2 82.4 3 - 183.4 - - - - - -
RAP [Cao et al.(2024)Cao, Tang, Huang, Jin, Zhang, Liu, Chen, Liang, Yuan, and Li] - ACL’24 44.8 71.4 81.5 - 14.4 197.7 - - - - - - 42.6 70.4 79.6 - 18.0 192.6
QB-Norm [Bogolin et al.(2022)Bogolin, Croitoru, Jin, Liu, and Albanie] QB CVPR’22 47.2 73.0 83.0 2 - 203.2 - - - - - - 43.3 71.4 80.8 2 - 195.5
DiCoSA [Jin et al.(2023b)Jin, Li, Cheng, Huang, Wang, Yuan, Liu, and Chen] QB IJCAI’23 47.5 74.7 83.8 2 13.2 206.0 - - - - - - 45.7 74.6 83.5 2 11.7 203.8
DiffusionRet [Jin et al.(2023c)Jin, Li, Cheng, Li, Ji, Liu, Yuan, and Chen] QB ICCV’23 48.9 75.2 83.1 2 12.1 207.2 - - - - - - 48.9 75.5 83.3 2 14.1 207.7
EMCL-Net [Jin et al.(2022)Jin, Huang, Liu, Wu, Ge, Song, Clifton, and Chen] DSL NeurIPS’22 51.6 78.1 85.3 1 - 215.0 - - - - - - - - - - - -
EMCL-Net [Jin et al.(2022)Jin, Huang, Liu, Wu, Ge, Song, Clifton, and Chen] DSL NeurIPS’22 - - - - - - - - - - - - 47.6 73.5 82.8 2 11.9 203.9
TS2-Net [Liu et al.(2022)Liu, Xiong, Xu, Cao, and Jin] DSL ECCV’22 51.1 76.9 85.6 1 11.7 213.6 - - - - - - 47.4 74.1 82.4 2 12.9 203.9
TABLE [Chen et al.(2023b)Chen, Wang, Lin, Qi, Ma, and Shan] DSL AAAI’23 52.3 78.4 85.2 1 11.4 215.9 - - - - - - 49.1 75.6 82.9 2 14.8 207.6
UATVR [Fang et al.(2023)Fang, Wu, Liu, Zhou, Song, Wang, Shu, Ji, and Wang] DSL ICCV’23 49.8 76.1 85.5 2 12.3 211.4 - - - - - - - - - - - -
RAP [Cao et al.(2024)Cao, Tang, Huang, Jin, Zhang, Liu, Chen, Liang, Yuan, and Li] DSL ACL’24 - - - - - - - - - - - - 47.1 74.1 82.4 - 13.9 203.6
TGIF Charades YouCook2
RIVRL [Dong et al.(2022b)Dong, Wang, Chen, Qu, Li, He, and Wang] - TCSVT’22 12.2 26.8 35.4 30 - 74.4 - - - - - - - - - - - -
X-Pool [Gorti et al.(2022)Gorti, Vouitsis, Ma, Golestan, Volkovs, Garg, and Yu] - CVPR’22 - - - - - - 16.1 35.2 44.9 14 67.2 96.2 - - - - - -
AME-Net [Han et al.(2022)Han, Chen, Zhang, Wang, and Chen] - TOMM’22 - - - - - - - - - - - - 7.6 21.5 32.8 28 - 61.9
EMCL-Net [Jin et al.(2022)Jin, Huang, Liu, Wu, Ge, Song, Clifton, and Chen] - NeurIPS’22 13.2 27.3 34.8 36 327.5 75.3 16.4 36.3 47.3 13 69.5 100.0 11.8 28.8 39.8 19 121.6 80.4
RoME [Satar et al.(2022)Satar, Zhu, Zhang, and Lim] - CoRR’22 - - - - - - - - - - - - 6.3 16.9 25.2 53 - 48.4
SwAMP [Kim(2023)] - AISTATS’23 - - - - - - - - - - - - 9.4 24.9 35.3 22 - 69.6
MuMUR [Madasu et al.(2023)Madasu, Aflalo, Stan, Rosenman, Tseng, Bertasius, and Lal] - Inf. Retr. J.’23 - - - - - - 16.6 37.5 50.0 10 52.7 104.1 - - - - - -
UCoFiA [Wang et al.(2023b)Wang, Sung, Cheng, Bertasius, and Bansal] - ICCV’23 13.0 27.1 34.7 35 323.8 74.8 16.4 37.2 47.7 12 74.6 101.3 12.4 30.4 41.0 17 123.5 83.8
LTME [Cheng et al.(2024a)Cheng, Kong, Wang, Qu, and Jiang] - ICASSP’24 10.6 24.1 32.2 - - 66.9 - - - - - - - - - - - -
BiC-Net [Han et al.(2024)Han, Zeng, Shi, Xiao, Chen, and Chen] - TOMM’24 - - - - - - - - - - - - 8.7 23.9 33.5 31 - 66.1
EMCL-Net [Jin et al.(2022)Jin, Huang, Liu, Wu, Ge, Song, Clifton, and Chen] DSL NeurIPS’22 14.8 29.4 36.6 31 317.3 80.8 18.2 38.2 49.3 11 63.1 105.7 12.6 31.7 42.5 17 104.5 86.8
Table 10: Comparison with SOTA across all the datasets. -: unreported results. our reproduced results. IS: Inference Strategy. The highest performance in each block is marked in bold.
Appendix E Full Tables of Experiments
Tab. 9 details the results presented in Fig. 4 and reports Recall at $L=1,5,10$ ($R@L$), median rank $MR$, mean rank ($MeanR$) and $RSum$ for all the datasets when using all the inference strategies. Similarly, Tab. 10 reports the full table of Tab. 3. Tab. 11 is the full table of the ablation Tab. 7 in the main paper. Tabs 11 to 14 shows the full tables with all the metrics of the ablation experiments in Sec. 4.3 of the main paper.
baseline 52.1 77.3 85.9 1 12.9 215.3
+VT 52.2 77.3 85.7 1 10.4 215.2
+$T$ 52.0 77.6 86.0 1 10.4 215.6
+VT+$T$ 53.2 77.7 85.3 1 10.0 216.2
Table 11: Ablation on architecture design.
Captions Blip2 PG 51.2 76.2 85.0 1 11.2 212.4
Captions Blip2 L3.1 51.6 76.7 85.5 1 10.9 213.8
Captions VL2 L3.1 50.4 75.9 84.7 1 11.5 211.0
Tags VL2 L3.1 53.2 77.7 85.3 1 10.0 216.2
Table 12: Ablation on auxiliary inputs. PG: PEGASUS. L3.1: Llama3.1. VL2: VideoLLaMA2.
Foundation Models
VT $T$ $R@1\uparrow$ $R@5\uparrow$ $R@10\uparrow$ $MR\downarrow$ $MeanR\downarrow$ $RSum\uparrow$
VL L2 52.0 77.5 84.9 1 10.4 214.4
VL2 L3.1 53.2 77.7 85.3 1 10.0 216.2
Table 13: Ablation on foundation models. VL: Video-LLaMA. VL2: VideoLLaMA2. L2: Llama2. L3.1: Llama3.1.
Inference Tags $R@1\uparrow$ $R@5\uparrow$ $R@10\uparrow$ $MR\downarrow$ $MeanR\downarrow$ $RSum\uparrow$
(i) original 53.2 77.7 85.3 1 10.0 216.2
(ii) original wrong 50.4 76.5 84.5 1 10.8 211.4
(iii) self-correct 53.9 78.1 86.1 1 9.4 218.1
(iv) self-correct wrong 49.9 76.0 84.8 2 11.1 210.7
Table 14: Ablation on quality of auxiliary tags.
Appendix F Additional Ablations
In this section, we ablate some additional design choices of MAC-VR. First, we ablate $K$ that controls the number of concepts/latent factors. We evaluate within the following range $K\in{4,8,16,32}$. In Tab. 16, the performance improves reaching the best result with $K=8$ and then decreases. This shows that a small number of concepts limits the ability to leverage fine-grained information, whereas a larger value reduces the dimensionality ofceach concept limiting the discriminability of the concept itself.